[figure] [figure] 30 December 1996 Monday 31 December 1996 Tuesday 1 January 1997 Wednesday

New Year 10:00 Wednesday call 16:00 Nancy 18:00–20:00 Dinner

2 January 1997 Thursday Friday 27 December 1996 Saturday 28 December 1996 Sunday 29 December 1996 [figure] [figure] 6 January 1997 Monday

Monday event 14:00–16:00 Monday meeting

7 January 1997 Tuesday 8 January 1997 Wednesday 9 January 1997 Thursday Friday 3 January 1997

Third of month 10:00–13:00 Koffee klatch

Saturday 4 January 1997 Sunday 5 January 1997